Find all numbers for which the rational expression is undefined
(n^3-8n)/(n^2-4)

Answer by nerdybill(7384)

It would only be undefined if the denominator is zero.
so, to find the values, we set (n^2-4) to zero and solve for n:
(n^2-4) = 0
n^2 = 4
n = sqrt(4)
n = {-2, 2} (THE ANSWER)